A small block with mass 0.221 kg is sliding down a frictionless ramp that is inclined at an angle of 53.1
Answer #1

a) normal force = gmcos(theta) = 9.8*0.221*cos(53.1) = 1.3004 N

b) force = 9.8*0.221*sin(53.1) = 1.732 N
Answer #2

normal force = mg cos theta

= 0.221 * 9.80 * cos53.1

= 1.3 N

part b)

F g = mg sin theta

= 0.221 * 9.8 * sin (53.1)

= 1.73 N